NOVEL APPROACHES TO DOPPLER-FREE ION IMAGING

\underline{CUNSHUN HUANG}, WEN LI, SRIDHAR A. LAHANKAR, MYUNG HWA KIM, BAILIN ZHANG, ARTHUR G. SUITS, Department of Chemistry, Wayne State University, Detroit, MI 48202.

We have recently developed novel approaches to Doppler-free imaging, named ”Two-color reduced Doppler ion imaging (TCRD)”^{a}. and ”Doppler-free/Doppler-sliced ion imaging”^{b}. These approaches offer advantages for application in a variety of situations. The TCRD approach is suitable for application in any system in which one would like to exploit a high power fixed frequency laser, or move a probe transition away from a region of background interference. It may be very useful in PHOFEX spectroscopy to encompass the entire Doppler profile of the detected species. The Doppler-free/ Doppler slicing approach is most suited for imaging studies and H or D atom detection. Each method improves the probe sensitivity and signal-to-noise ratio.

aC. Huang, W. Li, M. H. Kim, and A. G. Suits, J. Chem. Phys. 125, 121101 (2006).

bC. Huang, S. A. Lahankar, M. H. Kim, B. Zhang, and A. G. Suits, Phys. Chem. Chem. Phys. 8, 4652 (2006).